A propeller with three blades of mass 12 kg each , length of 1.25m each are 120 degrees relative to each other. If a torque N.m acts on this, what is the moment of inertia of the propeller? How long will it take to reach an angular velocity of 2000rad/s starting from rest?


1
Expert's answer
2021-12-03T12:44:44-0500

The moment of inertia of the propeller:


"I=mL^2."

Upon the torque "\tau", the angular acceleration is


"\alpha=\tau/I=\tau/(mL^2)=\tau/18.75\text{ rad/s}^2."


The time it takes to reach 2000 rad/s stating from rest:


"t=\frac{\omega}{\alpha}=\frac{\omega I}{\tau}=\frac{\omega mL^2}{\tau}=\frac{37500}\tau\text{ s}."